# On complexity of restricted fragments of Decision DNNF

### v1

@@ 3.3 Proofs of Theorem 9 and 10
Let $B$ be an obdd obeying an order $\pi$ and let ${\bf g}$ be an assignment to a prefix of $\pi$. Assume that $\emptyset\subset\mathcal{S}(B)|_{\bf g}\subset(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))^{\{0,1\}}$. Then $L({\bf g})$ is a singleton. Let $u({\bf g})$ be the only element of $L({\bf g})$. Then $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B_{u({\bf g})})\times X^{\{0,1\}}$, where $X=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\mathsf{var}(\mathcal{S}(B_{u({\bf g})})$.
@@
We observe that $|L({\bf g})|\leq 1$. Indeed, in case of two distinct nodes of $L({\bf g})$, we employ the argument as in the proof of the second statement of Lemma 2 to demonstrate that these nodes must have a $\wedge_{d}$ node as their common ancestor. As an obdd does not have such nodes, this possibility is ruled out. If we assume that $L({\bf g})=\emptyset$ then, taking into account that $\mathcal{S}(B)|_{\bf g}\neq\emptyset$, we conclude that the only node of $B[{\bf g}]$ besides the incomplete decision node is the $true$ sink. However, this means that all possible extensions of ${\bf g}$ are satisfying assignments of $B$ again contradicting our assumptions. We thus conclude that $L({\bf g})$ is indeed a singleton. The remaining part of the claim is immediate from Lemma 3. ∎
Note that Claim 3 formalizes the approach to proving obdd lower bounds that we outlined in Section 3.1. In particular, $u({\bf g})$ can be identified straight away without resorting to unbreakable sets. We thus proceed to define a fooling set $\mathcal{F}_{0}$ and to demonstrate that for every two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}_{0}$, $u({\bf g}_{1})\neq u({\bf g}_{2})$.
Let $\pi$ be an order obeyed by an obdd $B$ representing $\varphi(G)$ such that $M$ crosses $B$. Let $U,W\subseteq V(G)$ be two subsets of $V(G)$ such that $M$ is an induced matching of $G$ between $U$ and $W$ and each vertex of $U$ is ordered by $\pi$ before each vertex of $W$. Let $\pi_{0}$ be the prefix of $\pi$ whose last vertex is the last vertex of $U$ in $\pi$. Let $\mathcal{F}_{0}$ be the set of all assignments over $\pi_{0}$ that assign all variables of $\pi_{0}\setminus U$ by $1$ and assign at least one variable of $U$ by 00. It is not hard to observe that each ${\bf g}\in\mathcal{F}_{0}$ satisfies the premises of Claim 3.
@@
Let ${\bf g}_{1},{\bf g}_{2}$ be two distinct elements of $\mathcal{F}_{0}$ Then $u({\bf g}_{1})\neq u({\bf g}_{2})$.
@@
Assume the opposite. Let $U=\{u_{1},\dots,u_{q}\}$, $W=\{w_{1},\dots,w_{q}\}$ so that $M=\{\{u_{1},w_{1}\},\dots,\{u_{q},w_{q}\}\}$. Assume w.l.o.g. that there is $1\leq i\leq q$ such that ${\bf g}_{1}(u_{i})=0$ while ${\bf g}_{2}(u_{i})=1$.
Let ${\bf h}_{2}$ be an assignment over $V(G)$ such that ${\bf g}_{2}\subseteq{\bf h}_{2}$, ${\bf h}_{2}(w_{i})=0$ and ${\bf h}_{2}(w)=1$ for each other $w\in\pi\setminus\pi_{0}$. We observe that ${\bf h}_{2}$ satisfies $\varphi(G)$. Indeed, assume that ${\bf h}_{2}$ falsifies some $(u\vee v)$. By definition, both $u$ and $v$ belong to $V(M)$ but $\{u,v\}\in E(G)\setminus M$. But this is a contradiction to $M$ being an induced matching.
Let ${\bf h}=Proj({\bf h}_{2},\mathsf{var}(B_{u({\bf g}_{2})}))$. We note that $w_{i}\in\mathsf{var}({\bf h})$ and ${\bf h}(w_{i})=0$ simply by definition of ${\bf h}_{2}$. By Claim 3, ${\bf h}\in\mathcal{S}(B_{u({\bf g}_{2})})$ and hence, by our assumption, ${\bf h}\in\mathcal{S}(B_{u({\bf g}_{1})})$ implying that ${\bf g}_{1}\cup{\bf h}$ is a satisfying assignment of $\varphi(G)$. However, this is a contradiction since ${\bf g}_{1}\cup{\bf h}$ falsifies $(u_{i}\vee v_{i})$
It follows from Claim 4 that $|B|\geq|\mathcal{F}_{0}|$. It remains to notice that the projections of the elements of $\mathcal{F}_{0}|$ to $U$ include all possible assignments to $U$ but one. Hence, $|\mathcal{F}_{0}|\geq 2^{q}-1$.

### v2

@@ 4.2 Proof of Theorem 19
Let ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$ be a non-sink node and let ${\bf v}_{1}$ and ${\bf v}_{2}$ be the children of $\bf v$} in $B^{\prime}$. Then $em({\bf v}_{1})$ and $em({\bf v}_{2})$ are the children of $em({\bf v})$. Equivalently, if ${\bf v}_{1}$ and ${\bf v}_{2}$ are the children of $\bf v$} in $B^{\prime\prime}$ then $em^{-1}({\bf v}_{1})$ and $em^{-1}({\bf v}_{2})$ are the children of $\bf v$} in $B^{\prime}$.
The following statement is also immediate by construction.
Claim 4.
$\mathsf{var}(B^{\prime\prime}_{(u_{1},s_{1}+u)})=\mathsf{var}(B^{\prime}_{u})$.
Next, we observe that the $Var$ set remains the same for the subgraphs rooted by the nodes whose names are not changed by the transformation.
Claim 5.
Let ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$. Then $\mathsf{var}(B^{\prime}_{\bf v})=\mathsf{var}(B^{\prime\prime}_{\bf v})$
@@
By bottom up induction. The claim is trivially true for the sinks. Suppose that $\bf v$} is not a sink and let ${\bf v}_{1}$ and ${\bf v}_{2}$ be the children of $\bf v$} in $B^{\prime}$. By construction, if $\bf v$} is a variable node that the label of $\bf v$} is the same in $B^{\prime}$ and in $B^{\prime\prime}$. Hence the claim follows from the combination of the induction assumption and Claims 3 and 4. ∎
Now, we begin our reasoning towards establishing the embeddability.
@@
$B^{\prime\prime}_{(u_{1},s_{1}+u)}$ is a $\wedge_{d}$-fbdd embeddable into $\pi$.
@@
By construction and assumption about $B^{\prime}$, there are intervals $\pi_{1}$ and $\pi_{2}$ of $\pi$ such that such that $\pi_{1}$ precedes $\pi_{2}$ in $\pi$ and $B^{\prime}_{(u_{j},s_{j})}$ respects $\pi_{j}$ for each $j\in[2]$. is the same as being embeddable into $\pi$ according to Lemma 18
Claim 7.
For each ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$, $B^{\prime\prime}_{\bf v}$ is a $\wedge_{d}$-fbdd embeddable into $\pi$.
Proof.
The statement is trivially true for sinks so suppose that $\bf v$} with has children ${\bf v}_{1}$ and ${\bf v}_{2}$ in $B^{\prime\prime}$. It follows from the combination Claims 3 and 6 and the induction assumption that each $B^{\prime\prime}_{{\bf v}_{j}}$ is a $\wedge_{d}$-fbdd embeddable into $\pi$.
It follows from the combination of Claims 3, 4, and 5 that for each $j\in[2]$, $\mathsf{var}(B^{\prime}_{em^{-1}({\bf v}_{j})})=\mathsf{var}(B^{\prime\prime}_{{\bf v}_{j}})$. With this in mind, assume first that $\bf v$} is a decision node. Then its label $\mathsf{var}({\bf v})$ is not changed in $B^{\prime\prime}$. As in $B^{\prime}$, all of $\mathsf{var}(B^{\prime}_{em^{-1}({\bf v}_{j})})$ for each $j\in[2]$ are below $\mathsf{var}(B^{\prime}_{\bf v})$, the same dependency preserves for $B^{\prime\prime}$ for all $\mathsf{var}(B^{\prime\prime}_{{\bf v}_{j}})$ for each $j\in[2]$. Combined with the emeddability of as specified in the end of the previoys paragraph, we conclude that $B^{\prime\prime}_{\bf v}$ is embeddable in $\pi$.
It remains to assume that $\bf v$} is a $\wedge_{d}$-node. Assume w.l.o.g. that the edge $({\bf v},em^{-1}({\bf v}_{1}))$ is the top one in $B^{\prime}$. It follows that there are non-overlapping intervals $\pi_{1}$ and $\pi_{2}$ of $\pi$ with $\pi_{1}$ preceding $\pi_{2}$ and such that for each $j\in[2]$ $\mathsf{var}(B^{\prime}_{em^{-1}({\bf v}_{j})})\subseteq\pi_{j}$. We conclude that for each $j\in[2]$ $\mathsf{var}(B^{\prime\prime}_{{\bf v}_{j}})\subseteq\pi_{j}$ and hence $B^{\prime\prime}_{\bf v}$ is embeddable into $\pi$. ∎
Claim 8.
$B^{\prime\prime}_{source(B^{\prime\prime})}$ is a $\wedge_{d}$-fbdd embeddable into $\pi$.
Proof.
[…]
Claim 9.
$f(B^{\prime\prime}_{(u_{1},s_{1}+u)})=f(B^{\prime}_{u})$.
Proof.
Immediate by construction and from Lemmma 25. ∎
Claim 10.
For each ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$, $f(B^{\prime})_{\bf v}=f(B^{\prime\prime})_{\bf v}$
Proof.
We argue by bottom up induction. The statement is trivially true for sinks so suppose that $\bf v$} is a non-sink node of $B^{\prime\prime}$ with children ${\bf v}_{1}$ and ${\bf v}_{2}$. It follows from the combination Claims 3 and 9 and the induction assumption that for each $j\in[2]$ $f(B^{\prime\prime}_{{\bf v}_{j}})=f(B^{\prime}_{em^{-1}({\bf v}_{j})})$.
Assume that $\bf v$} is a variable node labelled with a variable $x$. Assume further that in $B^{\prime}$ the edge $({\bf v},em^{-1}({\bf v}_{1}))$ is labelled with $1$ and the edge $({\bf v},em^{-1}({\bf v}_{2}))$ is labelled with 0. Clearly, in $B^{\prime\prime}$, $\bf v$} remains a variable node labelled with $x$, $({\bf v},{\bf v}_{1})$ is labelled with $1$ and $({\bf v},{\bf v}_{2})$ is labelled with 0. Hence, the statement is immediate from the previous paragraph. If $\bf v$} is a $\wedge_{d}$-node in $B^{\prime}$ it remains so in $B^{\prime\prime}$ and hence the statement is again immediate from the previous paragraph. ∎
The following is immediate by construction
Claim 11.
$B^{\prime\prime}[V(B^{\prime})\cap V(B^{\prime\prime})]=B^{\prime}[V(B^{\prime})\cap V(B^{\prime\prime})]$, in particular the roles of the nodes, the labels on the nodes (if any) and the labels on the edges (if any) are all preserved.
We continue to observe that the roles of outgoing edges of $\wedge_{d}$ notes (the top and the bottom ones) are preserved.
Claim 12.
Let ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$ be a conjunction node. Let ${\bf w}_{1}$ be a child of $\bf v$} in $B^{\prime\prime}$. Then the role of $({\bf v},{\bf w}_{1})$ in $B^{\prime\prime}$ is the same as the one of $({\bf v},em^{-1}({\bf w}_{1}))$ in $B^{\prime}$
Proof.
Let ${\bf w}_{0}$ be the other child of $\bf v$} in $B^{\prime\prime}$. Due to the embeddability in $\pi$, there are non-overlapping intervals $\pi_{1}$ and $\pi_{0}$ of $\pi$ such that $\mathsf{var}(B^{\prime}_{em^{-1}({\bf w}_{1})})\subseteq\pi_{1}$ and $\mathsf{var}(B^{\prime}_{em^{-1}({\bf w}_{0})})\subseteq\pi_{0}$. The roles of the edges $({\bf v},em^{-1}({\bf w}_{1}))$ and $({\bf v},em^{-1}({\bf w}_{0}))$ are completely determined by the relative location of $\pi_{1}$ and $\pi_{0}$. It follows from the combination of Claims 3, 4, and 5 that $\mathsf{var}(B^{\prime}_{em^{-1}({\bf w}_{1})})=\mathsf{var}(B^{\prime\prime}_{{\bf w}_{1}})$ and $\mathsf{var}(B^{\prime}_{em^{-1}({\bf w}_{0})})=\mathsf{var}(B^{\prime\prime}_{{\bf w}_{0}})$. Hence the role of $({\bf v},{\bf w}_{1})$ is completely determined by the location of $\pi_{1}$ on $\pi_{0}$ relative to $\pi$ and the dependency is the same as in $B^{\prime}$. ∎
The first statement of the lemma follows from the combination of Claims 6 and 8, the second statement follows from the combination of Claims 9 and 10.
For the third statement, let $P$ be a path in $B^{\prime\prime}$ from $source(B^{\prime\prime})$ to $\bf v$} such that $ir_{\pi,B^{\prime\prime}}({\bf v})=ir_{\pi,B^{\prime\prime}}(P)$. If $P$ is also a path in $B^{\prime}$ then the statement is immediate from Claim 12. Otherwise, by Claim 11, $P$ must contain vertices of $D$. By construction, we can identify nodes ${\bf w}_{0}$ and ${\bf w}_{1}$ of $D$ so that $P=P_{1}+P_{2}+P_{3}$, where $P_{1}$ is the prefix of $P$ ending at ${\bf w}_{0}$, $P_{2}$ is the subpath of $P$ between ${\bf w}_{0}$ and ${\bf w}_{1}$, $P_{3}$ is the suffix of $P$ starting at ${\bf w}_{1}$ and $V(P)\cap V(D)=V(P_{2})$. We note that neither $P_{2}$ nor $P_{3}$ contain conjunction nodes. Hence, $ir_{\pi,B^{\prime\prime}}(P)=ir_{\pi,B^{\prime\prime}}(P_{1})$. Further on, by construction $B^{\prime}$ has a path $P^{\prime}=P^{\prime}_{1}+P^{\prime}_{3}$ where $P^{\prime}_{1}$ obtained from $P_{1}$ by replacing its last node with $u$ and $P^{\prime}_{3}$ is obtained from $P_{3}$ by replacing its first node with $u$. We observe that the outgoing edge of $u$ in $P^{\prime}$ is the bottom one and that $ir_{\pi,B^{\prime\prime}}(P_{1})=ir_{\pi,B^{\prime}}(P^{\prime}_{1})$ by Claim 12. We conclude that $ir_{\pi,B^{\prime}}(P^{\prime})=ir_{\pi,B^{\prime\prime}}(P)$. As $P^{\prime}$ is a path in $B^{\prime}$ from $source(B^{\prime})$ to $\bf v$}, this proves the third statement.
For the fourth statement, let $P$ be a path in $B^{\prime\prime}$ from $source(B^{\prime\prime})$ to $(u^{\prime},s^{\prime}+u)$ such that $ir_{\pi,B^{\prime\prime}}(P)=ir_{\pi,B^{\prime\prime}}((u^{\prime}s^{\prime}+u))$. By construction $P=P_{1}+P_{2}$ where $P_{1}$ is the prefix of $P$ ending at $(u_{1},s_{1}+u)$ and $P_{2}$ is the suffix of $P$ starting at $(u_{1},s_{1}+u)$. We note that $P_{2}$ does not contain conjunction nodes and hence $ir_{\pi,B^{\prime\prime}}(P)=ir_{\pi,B^{\prime\prime}}(P_{1})$. Further on, we observe that $P$ can be transformed by a path $P^{\prime}=P^{\prime}_{1}+P^{\prime}_{2}$ from $source(B^{\prime})$ where $P^{\prime}_{1}$ is obtained from $P_{1}$ by replacing the last node with $u$ and $P^{\prime}_{2}$ is obtained from $P_{2}$ by first replacing each node $(u^{*},s^{*}+u)$ with $(u^{*},s)$ and then adding $u$ at the beginning. We note that by Claim 12, $ir_{\pi,B^{\prime\prime}}(P_{1})=ir_{\pi,B^{\prime}}(P^{\prime}_{1})$. Furthermore, by construction, the outgoing edge of $u$ in $P^{\prime}_{2}$ is a top one, hence $ir_{\pi,B^{\prime}}(P^{\prime}_{2})\geq 1$. We conclude that $ir_{\pi,B^{\prime}}(P^{\prime})\geq ir_{\pi,B^{\prime\prime}}(P)+1$ thus establishing the fourth statement. ∎
Proof.
( of Theorem 19) Let $B^{*}$ be the output of $MakeOBDD(B,\pi)$. We are going to prove that $B^{*}$ is an ondd respecting $\pi$ of size at most $|B|^{ir_{\pi}(B)+1}$ and such that $f(B^{*})=f(B)$. As $B^{*}$ is obtained in a constructive way, the second statement of the theorem will be immediate.
It is clear from the description of the algorithm that the algorithm runs $q$ iterations where $q$ is at most the number of conjunction gates of $B$. The output of the algorithm is $B_{q}=B^{*}$. Applying the first two statements of lemma 26 inductively to each $B_{i}$, we conclude that $B_{q}$ is a $\wedge_{d}$-obdd embeddable into $\pi$ and $f(B_{q})=f(B)$. Since by construction, $B_{q}$ contains no conjunction gates, we conclude that $B^{*}$ is an $\wedge_{d}$-fbdd embeddable into $\pi$ and hence an obdd respecting $\pi$.
To establish an upper bound on the size of $B^{*}$, let us first replace $B_{i}$ with $i$ in the subscript for $ir$. That is, instead of $ir_{\pi,B_{i}}({\bf u})$, we write $ir_{\pi,i}({\bf u})$. Also for a decision node ${\bf u}=(u^{\prime},s^{\prime})$, let $len({\bf u})$ be the length of $s^{\prime}$. With this in mind, let us first prove the following claim.
Claim 13.
For each $0\leq i\leq q$ and a decision node ${\bf u}\in V(B_{i})$, $ir_{\pi,i}({\bf u})\leq ir_{\pi}(B)-len({\bf u})$.
Proof.
By induction on $i$. For each decision node ${\bf u}\in B_{0}$, $len({\bf u})=0$. Then, by Lemma 24, $ir_{\pi,0}({\bf u})\leq ir_{\pi}(B)=ir_{\pi}(B)-len({\bf u})$. Hence, the statement holds for $i=0$. Assume that $i>0$. Let ${\bf u}\in V(B_{i-1})\cap V(B_{i})$ be a decision node. By the third statement of Lemma 26 and the induction assumption, $ir_{\pi,i}({\bf u})\leq ir_{\pi,i-1}({\bf u})\leq ir_{\pi}(B)-len({\bf u})$. Let ${\bf u}\in V(B_{i})\setminus V(B_{i-1})$ be a decision node. By the fourth statement of Lemma 26, there is ${\bf u}_{0}\in{\bf B}_{i-1}$ such that $len({\bf u})=len({\bf u}_{0})+1$ and $ir_{\pi,i}({\bf u})\leq ir_{\pi,i-1}({\bf u}_{0})-1$. Hence, by the induction assumption, $ir_{\pi,i}({\bf u})\leq ir_{\pi}(B)-len({\bf u}_{0})-1=ir_{\pi}(B)-len({\bf u})$ as required. ∎
It follows that for each decision node ${\bf u}\in V(B_{q})$, $len({\bf u})\leq ir_{\pi}(B)$. Indeed, assume that this is not the case for some ${\bf u}\in V(B_{q})$. Then, by Claim 13, $ir_{\pi,i}({\bf u})\leq ir_{\pi}(B)-len({\bf u})<0$ in contradiction to the definition of the irregularity index of a node. We conclude that the names of the nodes are uniquely determined by sequences of at most $ir_{\pi}(B)+1$ elements of $B$, (the first coordinate plus a sequence of at most $ir_{\pi}(B)$ conjunction nodes). Clearly, the number of such distinct names is upper-bounded by $n^{ir_{\pi}(B)+1}$. ∎
4.3 Proof of Theorem 20
Let $f_{0}(s,x_{1},x_{2},y_{1},y_{2})$ be a Boolean function such that an assignment $\bf a$} to $\{s,x_{1},x_{2},y_{1},y_{2}\}$ is satisfying if and only if the following holds.
  1. 1.
    If ${\bf a}(s)=0$ then $\bf a$} does not map both $x_{1}$ ad $x_{2}$ to zeroes and $\bf a$} does not map both $y_{1}$ and $y_{2}$ to zeroes.
  2. 2.
    If ${\bf a}(s)=1$ then $\bf a$} does not map both $x_{1}$ ad $x_{2}$ to $1$. and $\bf a$} does not map both $y_{1}$ and $y_{2}$ to $1$.
For each $i\in\{0,1\}$, we denote by $D[x,y,i]$ an obdd querying variables $x$ and $y$ in the order listed. $f(D[x,y,i])$ is true on all the assignments but $\{(x,i),(y,i)\}$. Further on, let $D[s,x_{1},x_{2},y_{1},y_{2}]$ be defined as a $\wedge_{d}$-fbdd whose source $u$ is a variable node labelled with $s$. Let $u_{0}$ and $u_{1}$ be the children of $u$ such that $(u,u_{0})$ is labelled with 0 and $(u,u_{1})$ is labelled with $1$. Each $u_{i}$ is a $\wedge_{d}$ node whose children are $D[x_{1},x_{2},i]$ and $D[y_{1},y_{2},i]$. All the $True$ sinks of the latter two are contracted into a single $True$ sink and all the $False$ sinks are contracted into a single $False$ sink. We observe that
$f_{0}(s,x_{1},x_{2},y_{1},y_{2})=\bigvee_{i\in\{0,1\}}s^{i}\wedge D[x_{1},x_{2},i]\wedge D[y_{1},y_{2},i]$ (1)
where $s^{1}=s$ and $s^{0}=\neg s$. The next lemma follows from (1) and by construction.
Lemma 27.
$f(D[s,x_{1},x_{2},y_{1},y_{2}])=f_{0}(x_{1},x_{2},y_{1},x_{2})$
We now consider a variation $D^{-}[s,x_{1},x_{2},y_{1},y_{2}]$ of $D[s,x_{1},x_{2},y_{1},y_{2}]$ that differs from the former is that $True$ sinks are not all contacted together. In particular, we have the xTruex-True sink obtained by contracting $True$ sinks of graphs $D[x_{1},x_{2},i]$ and the yTruey-True sink obtained by contracting together $True$ sinks of the graphs $D[y_{1},y_{2},i]$.
We now define a $\wedge_{d}$-fbdd $B_{n}$ over variables $s_{1},x_{1,1},x_{1,2},y_{1,1},y_{1,2},\dots s_{n}.x_{n,1},x_{n,2},y_{n,1},y_{n,2}$. The building blocks for the graph are $D^{-}[s_{i},x_{i,1},x_{i,2},y_{i,1},y_{i,2}]$ for each $i\in[n-1]$ and $D[s_{n},x_{n,1},x_{n,2},y_{n,1},y_{n,2}]$. For each $i\in[n-1]$, we identify the xTruex-True sink of $D[s_{i},x_{i,1},x_{i,2},y_{i,1},y_{i,2}]$ with the source of $D[s_{i+1},x_{i+1,1},x_{i+1,2},y_{i+1,1},y_{i+1,2}]$. The remaining $True$ sinks are contacted into a single $True$ sink and the remaining $False$ sinks are contracted into a single $False$ sink.
In order to proceed, we need to upgrade our notation. We denote the set $\{s_{1},\dots,s_{n}\}$ by $S_{n}$ and the set $\{s_{i},\dots,s_{n}\}$ by $S^{i}_{n}$. Further on, we denote the sets $\{x_{1,1},x_{1,2},\dots,x_{n,1},x_{n,2}\}$ and $\{y_{1,1},y_{1,2},\dots,y_{n,1},y_{n,2}\}$ by $X_{n}$ and $Y_{n}$ respectively and the sets $\{x_{i,1},x_{i,2},\dots,x_{n,1},x_{n,2}\}$ and $\{y_{i,1},y_{i,2},\dots,y_{n,1},y_{n,2}\}$ by $X^{i}_{n}$ and $Y^{i}_{n}$ respectively. Next $\wedge_{i\in[n]}f_{0}(s_{i},x_{i,1},x_{i,2},y_{i,1},y_{i,2})$ is denoted by $f_{n}(S_{n},X_{n},Y_{n})$. Finally, $\wedge_{j\in\{i,\dots,n\}}f_{0}(s_{j},x_{j,1},x_{j,2},y_{j,1},y_{j,2})$ is denoted by $f^{i}_{n}(S^{i}_{n},X^{i}_{n},Y^{i}_{n})$.
Lemma 28.
$f(B_{n})=f_{n}(S_{n},X_{n},Y_{n})$.
Proof.
Let $u[1],\dots,u[n]$ be the respective sources of
$D^{-}[s_{1},x_{1,1},x_{1,2},y_{1,1},y_{1,2}],\dots D^{-}[s_{n-1},x_{n-1,1},x_{n-1,2},y_{n-1,1},y_{n-1,2}]$ and $D[s_{n},x_{n,1},x_{n,2},y_{n,1},y_{n,2}]$.
We prove by induction on $i$ going from $n$ down to $1$ that $f([B_{n}]_{u[i]})=f^{i}_{n}(S^{i}_{n},X^{i}_{n},Y^{i}_{n})$. Since $B_{n}]_{u[1]}=B_{n}$, this will imply the lemma. For $i=n$, this is just Lemma 27. So, we assume that $i<n$.
For $a\in\{0,1\}$, let $w_{a}$ be the source of $D[x_{i,1},x_{i,2},a]$. It is immediate from the induction assumption that
$f([B_{n}]_{w_{a}})=f(D[x_{i,1},x_{i,2},a])\wedge f^{i+1}_{n}(S^{i+1},X^{i+1}_{n},Y^{i+1}_{n})$ (2)
Further on, for each $a\in\{0,1\}$, let $v_{a}$ be the child of $u[i]$ such that $(u[i],v_{a})$ is labelled with $a$. It follows by construction and from (2) that
$f([B_{n}]_{v_{a}})=f(D[x_{i,1},x_{i_{2}},a])\wedge f(D[y_{i,1},y_{i,2},a])\wedge f^{i+1}_{n}(S^{i+1}_{n},X^{i+1}_{n},Y^{i+1}_{n})$ (3)
It then follows from (3) and by construction that
$f([B_{n}]_{u[i]}=(\bigvee_{a\in\{0,1\}}s_{i}^{a}\wedge f(D[x_{i,1},x_{i,2},a])\wedge f(D[y_{i,1},y_{i,2},a]))\wedge f^{i+1}_{n}(S^{i+1}_{n},X^{i+1}_{n},Y^{i+1}_{n})$ (4)
Combining (1), (3), and (4) we observe that
$f([B_{n}]_{u[i]}=f_{0}(s_{i},x_{i,1},x_{i,2},y_{i,1},y_{i,2})\wedge f^{i+1}_{n}(S^{i+1}_{n},X^{i+1}_{n},Y^{i+1}_{n})$ (5)
It only remains to observe that $f_{0}(s_{i},x_{i,1},x_{i,2},y_{i,1},y_{i,2})\wedge f^{i+1}_{n}(S^{i+1}_{n},X^{i+1}_{n},Y^{i+1}_{n})=f^{i}_{n}(S^{i}_{n},X^{i}_{n},Y^{i}_{n})$. ∎
In order to proceed, we need an easily verifiable lemma addressing the situation where an obdd is placed on top of $\wedge_{d}$-fbdd.
Lemma 29.
Let $\pi$ be a linear order. Let $B_{1}$ be an obdd with two sinks respecting $\pi$. Further on, let $B_{2}$ be a $\wedge_{d}$-fbdd embeddable into $\pi$. Assume that all of $\mathsf{var}(B_{1})$ occur in $\pi$ before all of $\mathsf{var}(B_{2})$. Let $B$ be a $\wedge_{d}$-fbdd obtained by identifying the positive sink of $B_{1}$ with the source of $B_{2}$ and contracting the negative sinks of $B_{1}$ and $B_{2}$. Then $B$ is embeddable into $\pi$ with $ir_{\pi}(B)=ir_{\pi}(B_{2})$.
Let $\pi_{n}=s_{1},x_{1,1},x_{1,2}\dots s_{n},x_{n,1},x_{n,2},y_{n,1},y_{n,2},\dots y_{1,1},y_{1,2}$ and let $\pi^{i}_{n}=s_{i},x_{i,1},x_{i,2}\dots s_{n},x_{n,1},x_{n,2},y_{n,1},y_{n,2},\dots y_{i,1},y_{i,2}$
Lemma 30.
$B_{n}$ is embeddable in $\pi_{n}$ with irregularity index $n$.
Proof.
Let $u[1],\dots,u[n]$ be as in the proof of Lemma 28. For each $i$ from $n$ down to $1$ we prove that $B_{n}]_{u[i]}$ is embeddable in $\pi_{n}$ with the irregularity index ni+1n-i+1. The lemma will then follow as $B_{n}]_{u[1]}=B_{n}$.
Assume first that $i=n$. For each $a\in\{0,1\}$, let $w_{x,a}$ be the source of $D[x_{n,1},x_{n,2},a]$ and let $w_{y,a}$ be the source of $D[y_{n,1},y_{n,2},a]$. We observe that since $(x_{n,1},x_{n,2})$ and $(y_{n,1},y_{n,2})$ are both intervals of $\pi_{n}$, both $B_{n}]_{w_{x,a}}$ and $B_{n}]_{w_{y,a}}$ are embeddable into $\pi_{n}$ with irregularity index 0 by Lemma 18. Next, let $v_{a}$ be the $\wedge_{d}$-node whose children are $w_{x,a}$ and $w_{y,a}$. Since $(x_{n,1},x_{n,2})$ and $(y_{n,1},y_{n,2})$ are disjoint intervals of $\pi_{n}$, we conclude that $B_{n}]_{v_{a}}$ is embeddable into $\pi_{n}$ with irregularity index $1$. We node that $\mathsf{var}([B_{n}]_{v_{a}})$ form a subinterval of $\pi_{n}$ occurring after $s_{n}$ hence a subinterval of $\pi_{n}^{>s_{n}}$. We conclude that, by construction $B_{n}]_{u[n]}$ is embeddable into $\pi_{n}$ with irregularity index $1$.
Assume now that $i<n$. We observe first that the order of $\mathsf{var}([B_{n}]_{u[i+1]})]$ under $\pi$ is $\pi_{n}^{i+1}$, clearly an interval of $\pi_{n}$. For each $a\in\{0,1\}$, we now reuse the notation for $w_{x,a}$ to the source of $D[x_{i,1},x_{i,2},a]$ and let $w_{y,a}$ be the source of $D[y_{i,1},y_{i,2},a]$. Since $(x_{i,1},x_{i,2})$ occurs in $\pi_{n}$ before $\pi^{i+1}_{n}$ we argue by construction, the induction assumption and Lemma 29 that $B_{n}]_{w_{x,a}}$ is embeddable into $\pi_{n}$ with irregularity index nin-i. Further on, we reuse $v_{a}$ to be the $\wedge_{d}$ node connecting $w_{x,a}$ and $w_{y,a}$ . We observe that since $(y_{i,1},y_{i,2})$ occurs in $\pi_{n}$ after $(x_{i,1},x_{i_{2}})+\pi_{i+1}$ $B_{n}]_{v_{a}}$ is embeddable into $\pi_{n}$ with irregularity index ni+1n-i+1. We further note that $Var([B_{n}]_{v_{a}})$ form an interval of $\pi_{n}$ occurring after $s_{i}$ and hence an interval of $\pi^{>s_{i}}_{n}$. Therefore, by construction $B_{n}]_{u[i]}$ is embeddable into $\pi_{n}$ with the irregularity index ni+1n-i+1 as required. ∎
Lemma 31.
Let $B$ be an obdd representing $f_{n}=f_{n}(S_{n},X_{n},Y_{n})$ and respecting $\pi_{n}$. Then $|B|\geq 2^{n}$.
Proof.
Let $\bf S$} be the set of all the assignments to $S_{n}\cup X_{n}$. that map each $x_{i,1}$ to 0 and each $x_{i,2}$ to $1$.
Claim 14.
For any two ${\bf a}_{1},{\bf a}_{2}\in{\bf S}$, $f_{n}|_{{\bf a}_{1}}\neq f_{n}|_{{\bf a}_{2}}$.
Since $\pi_{n}$ places all the variables in $\bf S$} before all the variables of $\mathsf{var}(f_{n})\setminus{\bf S}$, it follows from the claim and a well known facts about obdds that $|B|\geq|{\bf S}|$. By construction, $|{\bf S}|=2^{n}$ so the lemma follows. It this remains to prove the claim.
Since ${\bf a}_{1}\neq{\bf a}_{2}$, there is $i\in[n]$ such that say ${\bf a}_{1}(s_{i})=0$ whereas ${\bf a}_{1}(s_{1})=1$. It follows that there is no satisfying assignment for $f_{n}|_{{\bf a}_{1}}$ that assigns both $y_{i,1}$ and $y_{i,2}$ with zeroes while there is such an assignment for $f_{n}|_{{\bf a}_{2}}$. ∎
Proof.
(Of Theorem 20) For each $n$ we consider the function $f_{n}$ and the order $\pi_{n}$ of its variables. We note that, by construction, $B_{n}$ is of size linear in $n$. Furthermore, it follows from combination of Lemmas 30 and 31 that $f(B_{n})=f_{n}$ and $B_{n}$ respects $\pi_{n}$ with irregularity index $n$. On the other hand, it follows from Lemma 31 that if $f_{n}$ is represented by an obdd respecting $\pi_{n}$ then such an obdd must be of size at least $2^{n}$. ∎